please help me to solve this problems
1)9/10x+3=12
2)x/5-3=1
3)2x-1/3=5/3
4)6/x=2/7

Answer by ikdeep(189)

is your problems look like this

1)(9/10)(x+3)=12
2)(x/5)-3=1
3)(2x-1)/3=5/3
4)(6/x)=2/7.......
Please always use brackets when you give problems in computer as problem see differently on paper and on computer.
Please contact again with clarification......